If the above grandfather clock were put on a rocket ship that launched vertically with an acceleration of .50 g’s, what would be the time interval between ticks?
Answer #1

as we know time period os simple pendulum is

T = 2 pi sqrt ( L/g)

for constant L

T2 /T1 = sqrt ( g2/g1)

T2 /1 = sqrt ( 0.5 g/g)

T2 = 0.707 s

=========

Comment before rate in case any doubt, will reply for sure.. goodluck